# Perm: A Parametric Representation for Multi-Style 3D Hair Modeling

### v5

@@ Abstract
We present Perm, a learned parametric representation of human 3D hair designed to facilitate various hair-related applications. Unlike previous work that jointly models the global hair structure and local curl patterns, we propose to disentangle them using a PCA-based strand representation in the frequency domain, thereby allowing more precise editing and output control. Specifically, we leverage our strand representation to fit and decompose hair geometry textures into low- to high-frequency hair structures, termed guide textures and residual textures, respectively. These decomposed textures are later parameterized with different generative models, emulating common stages in the hair grooming process. We conduct extensive experiments to validate the architecture design of Perm, and finally deploy the trained model as a generic prior to solve task-agnostic problems, further showcasing its flexibility and superiority in tasks such as single-view hair reconstruction, hairstyle editing, and hair-conditioned image generation. More details can be found on our project page: https://cs.yale.edu/homes/che/projects/perm/.
@@ 4.2 Full Hair Representation
Figure 7: Comparison with GroomGen (Zhou et al., 2023) on random hairstyle synthesis. Full results are available in Fig. 27.
The most relevant previous work is GroomGen (Zhou et al., 2023) that learns hierarchical representations of 3D hair. As there is no publicly available GroomGen code, we implemented and trained it on the same augmented USC-HairSalon dataset. We also contacted the authors to obtain part of their official checkpoints to verify the correctness of our implementation. Detailed verification can be found in Appendix C.4, where we compared each network module with the available official checkpoints to demonstrate the correctness of our implementation. Through experiments with our implementation, we found the neural upsampler module in GroomGen, which is a GAN architecture, is very unstable to train, frequently leading to collapsed results. In Fig. 7 we visualize randomly synthesized hairstyles by sampling our parameter space and GroomGen’s latent space with the same Gaussian noise, demonstrating their collapsed output with weird shapes and curls. We further conducted a quantitative evaluation of reconstruction errors on the $17$ testing hairstyles, where our model achieves a position error of $1.658$ and curvature error of $0.769$, both of which are lower than GroomGen’s corresponding values (position error $2.570$, curvature error $6.199$). During quantitative evaluation, we also found that $100$ points per strand is insufficient to represent kinky hairstyles, and GroomGen’s output failed to capture the detailed curl patterns. Examples are provided in Fig. 25.
@@ C.4 Analysis of GroomGen
Refer to caption Refer to caption Refer to caption Refer to caption
GroomGen Reference Original Kinky Hair Downsampled Hair Reconstructed Hair
@@
Figure 25: GroomGen verification on a manually made kinky hairstyle.
To assess whether our implementation of GroomGen can reproduce kinky hairstyles similar to those shown in the GroomGen teaser (Fig. 25 $1$st column), we manually created a similar kinky hairstyle (Fig. 25 $2$nd column). We found that each strand required $200$ points to accurately capture the fine curls. However, GroomGen claims to predict $100$ points for each strand in its architecture. When we downsampled the strands to $100$ points, though the overall shape of the hairstyle remained intact, its curls were significantly degraded (Fig. 25 $3$rd column). We then searched for the optimal latent code to reconstruct the downsampled hair, yielding the results shown on the last column of Fig. 25. Despite these efforts, we found that the architecture struggled to faithfully reconstruct the hairstyle with accurate fine details.
@@ D.1 Random Hairstyle Synthesis
Figure 27: Full hair models synthesized from different Gaussian noise, with comparison to GroomGen (Zhou et al., 2023) trained on the same dataset. Our model and GroomGen are sampled using the same Gaussian noise. Hair colors are manually assigned for aesthetic purposes.
@@ Appendix E Limitations
Our work still suffers from several limitations. First, our training data is limited, as high-quality 3D hair assets are notoriously challenging to acquire. Consequently, certain intricate hairstyles, such as afro styles and various braided patterns, are underrepresented, which limits the generative capabilities of our model and leads to unnatural outputs (see Fig. 34). Addressing this issue requires the development of a systematic framework for efficiently capturing a diverse range of real-world data – a critical challenge that remains unresolved in the hair modeling area. However, Perm could serve as a pre-trained prior for efficient data capture, since it has the potential to fill in invisible parts of the hair, such as interiors and occluded regions, thereby accommodating more sparse image inputs. The captured hair data could also be used to fine-tune Perm, further bridging the domain gap between synthetic and real-world hairstyles. Second, our current single-view hair reconstruction pipeline adopts a relatively straightforward approach, relying on an optimization process that primarily focuses on per-pixel strand information encoded in the 2D rendered images. These supervisions may fail to capture fine strand-level details, yielding suboptimal outputs in extreme cases (see Fig. 35). To enhance reconstruction quality, existing methods (Chai et al., 2012; Nam et al., 2019) could be incorporated to trace continuous strand segments from single-view or multi-view input images, providing a more reliable geometric clue for our optimization. Additionally, more advanced optimization techniques could be explored. For example, integrating patch-based losses, such as LPIPS (Zhang et al., 2018), may help encourage the preservation of strand details at the patch level, thereby improving the fidelity of the reconstructed hairstyles.

### v6

@@ Abstract
We present Perm, a learned parametric representation of 3D human hair designed to facilitate various hair-related applications. Unlike previous work that jointly models the global hair structure and local curl patterns, we propose to disentangle them using a PCA-based strand representation in the frequency domain, thereby allowing more precise editing and output control. Specifically, we leverage our strand representation to fit and decompose hair geometry textures into low- to high-frequency hair structures, termed guide textures and residual textures, respectively. These decomposed textures are later parameterized with different generative models, emulating common stages in the hair grooming process. We conduct extensive experiments to validate the architecture design of Perm, and finally deploy the trained model as a generic prior to solve task-agnostic problems, further showcasing its flexibility and superiority in tasks such as single-view hair reconstruction, hairstyle editing, and hair-conditioned image generation. More details can be found on our project page: https://cs.yale.edu/homes/che/projects/perm/.
@@ 4.2 Full Hair Representation
Figure 7: Comparison with our implementation of GroomGen (Zhou et al., 2023) on random hairstyle synthesis. Full results are available in Fig. 27.
The most relevant previous work is GroomGen (Zhou et al., 2023) that learns hierarchical representations of 3D hair. As there is no publicly available GroomGen code, we implemented and trained it on the same augmented USC-HairSalon dataset. We also contacted the authors to obtain part of their official checkpoints to verify the correctness of our implementation. Detailed verification can be found in Appendix C.4, where we compared each network module with the available official checkpoints to demonstrate the correctness of our implementation. Through experiments with our implementation, we found the neural upsampler module in GroomGen, which is a GAN architecture, is very unstable to train, frequently leading to collapsed results. In Fig. 7 we visualize randomly synthesized hairstyles by sampling our parameter space and GroomGen’s latent space with the same Gaussian noise, demonstrating their collapsed output with weird shapes and curls. We further conducted a quantitative evaluation of reconstruction errors on the $17$ testing hairstyles, where our model achieves a position error of $1.658$ and curvature error of $0.769$, both of which are lower than GroomGen’s corresponding values (position error $2.570$, curvature error $6.199$). During quantitative evaluation, we also found that $100$ points per strand are insufficient to faithfully represent long kinky hairstyles. Examples are provided in Fig. 25.
@@ C.4 Analysis of GroomGen
Refer to caption Refer to caption Refer to caption
Original Kinky Hair Downsampled Hair Reconstructed Hair
@@
Figure 25: Verification of our GroomGen implementation on a manually made kinky hairstyle.
To assess whether our implementation of GroomGen can reproduce long kinky hairstyles, we manually created a test hairstyle (Fig. 25 $1$st column) which originally contains $200$ points per strand to accurately capture the fine curls. We downsampled the strands to $100$ points to fit our design, which led to a noticeable degradation in curl fidelity (Fig. 25 $2$nd column). We then searched for the optimal latent code to reconstruct the downsampled hair, yielding the results shown in the last column of Fig. 25. Despite these efforts, our implementation struggled to faithfully reconstruct the hairstyle with accurate fine details.
@@ D.1 Random Hairstyle Synthesis
Figure 27: Full hair models synthesized from different Gaussian noise, with comparison to our implementation of GroomGen (Zhou et al., 2023) trained on the same dataset. Our model and GroomGen are sampled using the same Gaussian noise. Hair colors are manually assigned for aesthetic purposes.
@@ Appendix E Limitations
Our work still suffers from several limitations. First, our training data is limited, as high-quality 3D hair assets are notoriously challenging to acquire. Consequently, certain intricate hairstyles, such as afro styles and various braided patterns, are underrepresented, which limits the generative capabilities of our model and leads to unnatural outputs (see Fig. 34). Addressing this issue requires the development of a systematic framework for efficiently capturing a diverse range of real-world data – a critical challenge that remains unresolved in the hair modeling area. However, Perm could serve as a pre-trained prior for efficient data capture, since it has the potential to fill in invisible parts of the hair, such as interiors and occluded regions, thereby accommodating more sparse image inputs. The captured hair data could also be used to fine-tune Perm, further bridging the domain gap between synthetic and real-world hairstyles.
@@
Second, our current single-view hair reconstruction pipeline adopts a relatively straightforward approach, relying on an optimization process that primarily focuses on per-pixel strand information encoded in the 2D rendered images. These supervisions may fail to capture fine strand-level details, yielding suboptimal outputs in extreme cases (see Fig. 35). To enhance reconstruction quality, existing methods (Chai et al., 2012; Nam et al., 2019) could be incorporated to trace continuous strand segments from single-view or multi-view input images, providing a more reliable geometric clue for our optimization. Additionally, more advanced optimization techniques could be explored. For example, integrating patch-based losses, such as LPIPS (Zhang et al., 2018), may help encourage the preservation of strand details at the patch level, thereby improving the fidelity of the reconstructed hairstyles.